Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
casanueva, mantovani, bonnand, gouaty - 8:57 Tuesday 22 January 2019 (44469) Print this report
Recovery of the morning

Yesterday morning we have started the recovery by aligning the arm cavities, without major problems. Then we encountered problems locking the PRC, until we realized that the phase was not well tuned. So we retuned both the phase of the B2 longitudinal PD and the quadrant photodiodes. Then everything went well until DF where we also had to retune the phase. Finally we reached LN1 and we tried to retune the phase of the final alignment as well.

We encountered problems closing the galvos, so both ROmains helped us by centering the beam on B1p and B5 quadrants. Also they put in the nominal position the SRM, which could have a big impact on the state of the quadrants. However this did not improved the situation. I managed to close all the loops but not in a robust way and with the BS gain divided by 2. We also saw that the vertical loops PR/BS were quite insensitive to the lines we injected, as if they would not see the corresponding DOF. ALso the working point of the PR would change from one lock to the other.

At the end of the shift we let romain align the OMCs

Comments to this report:
gouaty - 13:28 Tuesday 22 January 2019 (44477) Print this report

This is Michal who realigned the OMC, using the single bounce beam from NI aligned.

Search
×

Warning

×